OpenStudy (anonymous):

lim x->3 where x^2+5 does not equal 3 and 7 equals 3

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3} \left[\begin{matrix}x^2+5 & x \neq3 \\ 7 & x=3\end{matrix}\right]\]

OpenStudy (anonymous):

The limit is nine, but the value of the function is seven. Limit exists, because the limit from the left and right are the same, but it isn't equal to the function value, because of a discontinuity of the function at that point.

OpenStudy (anonymous):

wait how did you get 9 and get that the limt on the left and right are the same?

OpenStudy (anonymous):

the limit is 14. It doesn't matter if the function is defined or not at x = 3.

OpenStudy (anonymous):

Yeah....it really is bedtime....LOL Sorry about that, and good night.

OpenStudy (anonymous):

question?

OpenStudy (anonymous):

nope all good. thank you